# Effort: Efficient Orthogonal Modeling for Generalizable AI-Generated Image Detection

### v1

@@ 5 Conclusion
In this paper, we provide a new perspective to explain the failure reason of the generalization in AIGI detection, namely the asymmetry phenomena, where a naively trained detector very quickly shortcuts to the seen fake patterns, dominating the whole feature space and shaping it to be highly low-ranked, thereby limiting its expressivity and generalization. The key idea of this paper is to incorporate the rich semantic knowledge within the vision foundation models (VFMs) to expand the discrimination and feature space to higher-ranked, alleviating the model’s overfitting to seen fake patterns. A straightforward strategy is to directly fine-tune the VFM, but it risks distorting the learned semantic space, pushing the model back toward overfitting. To this end, we propose to decompose the original semantic space into two orthogonal subspaces for preserving semantic knowledge while learning forgery. We conduct extensive experiments on both deepfake and synthetic image detection benchmarks, showing the superior advantages in both generalization and efficiency.
Appendix
This appendix material provides additional experimental results, visualizations, and future works:
  • Algorithm procedure of the proposed approach (see Sec. A).
  • Additional experimental results and ablations (see Sec. B).
  • Additional visualizations and analysis (see Sec. D).
  • Future works and limitation discussion (see Sec. E).
Appendix A Algorithm Procedure Illustration
@@
Appendix B Additional Results and Ablations
@@
B.1 Results on GenImage Benchmark
In our manuscript, we present the benchmarking outcomes of the UniversalFakeDetect Dataset. Additionally, we report the results obtained from another widely utilized benchmark known as GenImage [103]. This GenImage dataset predominantly utilizes the Diffusion model for image generation, incorporating models such as Midjourney [56], SDv1.4 [68], SDv1.5 [68], ADM [18], GLIDE [59], Wukong [1], VQDM [22], and BigGAN [3]. Following the settings defined for GenImage, we designate SDv1.4 as the training set and the remaining models as the test set. Given the diverse image sizes within the GenImage dataset, images with a size smaller than 224 pixels are duplicated and subsequently cropped to 224 pixels, following [79]. We employ the same setting to re-implement FreqNet, FatFormer, and NPR, and also report the results of UnivFD and DRCT from [6].
@@ B.1 Results on GenImage Benchmark
ResNet-50 [27] CVPR 2016 54.9 99.9 99.7 53.5 61.9 98.2 56.6 52.0 72.1
DeiT-S [82] ICML 2021 55.6 99.9 99.8 49.8 58.1 98.9 56.9 53.5 71.6
Swin-T [49] ICCV 2021 62.1 99.9 99.8 49.8 67.6 99.1 62.3 57.6 74.8
CNNSpot [86] CVPR 2020 52.8 96.3 95.9 50.1 39.8 78.6 53.4 46.8 64.2
Spec [98] WIFS 2019 52.0 99.4 99.2 49.7 49.8 94.8 55.6 49.8 68.8
F3Net [64] ECCV 2020 50.1 99.9 99.9 49.9 50.0 99.9 49.9 49.9 68.7
GramNet [50] CVPR 2020 54.2 99.2 99.1 50.3 54.6 98.9 50.8 51.7 69.9
UnivFD [61] CVPR 2023 91.5 96.4 96.1 58.1 73.4 94.5 67.8 57.7 79.5
NPR  [79] CVPR 2024 81.0 98.2 97.9 76.9 89.8 96.9 84.1 84.2 88.6
FreqNet  [78] AAAI 2024 89.6 98.8 98.6 66.8 86.5 97.3 75.8 81.4 86.8
FatFormer [48] CVPR 2024 92.7 100.0 99.9 75.9 88.0 99.9 98.8 55.8 88.9
DRCT [6] ICML 2024 91.5 95.0 94.4 79.4 89.2 94.7 90.0 81.7 89.5
Ours 82.4 99.8 99.8 78.7 93.3 97.4 91.7 77.6 91.1
Table 7: Benchmarking results of cross-method evaluations in terms of Acc performance on the Genimage dataset. We follow [103] and use the SDv1.4 as the training set while others as the testing sets. We directly cite the results of ResNet-50, DeiT-S, Swin-T, CNNSpot, Spec, F3Net, and GramNet from [103]. We obtain the results of UnivFD and DRCT from [6], and FreqNet, NPR, and FatFormer by using the official checkpoints for reproduction. We report the Accuracy metric for comparison following [6].
@@
LipForensics [25] CVPR 2021 0.824 0.735 0.780
FTCN [100] ICCV 2021 0.869 0.740 0.805
HCIL [23] ECCV 2022 0.790 0.692 0.741
RealForensics [26] CVPR 2022 0.857 0.759 0.808
LTTD [24] NeurIPS 2022 0.893 0.804 0.849
AltFreezing [88] CVPR 2023 0.895
TALL-Swin [90] ICCV 2023 0.908 0.768 0.838
StyleDFD [12] CVPR 2024 0.890
NACO [97] ECCV 2024 0.895 0.767 0.831
@@
B.2 Comparison with Existing Video Detectors
@@
In the manuscript, we mainly compare our method with image detectors. Here, we provide an individual result to compare our approach with existing SOTA video detectors. Following [97, 90], we conduct evaluations on the widely-used CDF-v2 [45] and DFDC [15] using the video-level AUC metric. We have considered both the classical detectors such as LipForensics and the latest SOTA detectors such as NACO (ECCV’24) for a comprehensive comparison. Results in Tab. 8 demonstrate that our image-based approach achieves higher generalization performance in both CDF-v2 and DFDC, improving 4.5% and 3.9% points than the second-best video-based models. This further validates the superior generalization performance of our approach.
Methods CDF-v2 DFD DFDC Avg.
LoRA [37] 0.838 0.834 0.717 0.796
MoE-LoRA [38] 0.867 0.904
Dual-Adapter [71] 0.717 0.727
Ours 0.901 0.923 0.798 0.874
Table 9: Cross-dataset generalization evaluations with other adapter-based partially fine-tuned detectors. The results are cited from their original papers. The metric is frame-level AUC.
Archs. $n-r$ $r$ mAcc
UniFD [61] (Linear-Prob) 81.02
Baseline (FFT) 86.22
Ours Variant-1 256 92.13
Variant-2 64 93.68
Variant-3 16 94.45
Variant-4 4 94.37
Variant-5 1 95.19
LoRA Variant-1 256 91.42
Variant-2 64 91.06
Variant-3 16 91.89
Variant-4 4 93.53
Variant-5 1 93.03
Table 10: Ablation studies on synthetic image detection (UniversalFakeDetect Dataset) regarding the tunable $n-r$ values in SVD (Ours) and $r$ values in LoRA. All models are trained on ProGAN’s images and tested on 19 different kinds of generative models’ images. “FFT” indicates the full fine-tuning. “Linear-Prob” indicates fine-tuning FC layer only, where we reproduce the results from UniFD [61].
B.3 Within-Domain Evaluation Results
In the previous contents of the manuscript, we mainly show the cross-domain generalization results. For a more comprehensive evaluation, we provide our method’s within-domain evaluations. Since we use FF++ [70] (c23) for training, following DeepfakeBench [92], we train our model on FF++ (c23) and test it on FF++ (c23) and FF++ (c40), where c40 is a compression version of c23. We compare the results of our approach with other competing baseline models such as CNN-Aug, F3Net, SPSL, SRM, and RECCE from DeepfakeBench. We also compare our approach with the latest SOTA detector, i.e., CDFA [46] (ECCV’24) for a comprehensive comparison. Results in Tab. 11 show that generally, all detectors perform very well on FF++ (c23), as the training and testing distributions are similar. However, the performance largely drops when applying compression (tested on the c40 version). Also, our approach demonstrates the best results on FF++ (c40), showing the robustness of our detector. Notably, it can be seen that although CDFA achieves superior generalization results on cross-domain evaluation scenarios, it’s performance on within-domain evaluation is limited. In contrast, our approach shows high results on both within-domain and cross-domain evaluations. Regarding the model’s robustness toward the post-processing, we provide another robustness experiment following [32] (refer to Sec. B.5 for details).
B.4 Comparison with Existing Adapter-Based Detectors
Parameter-efficient fine-tuning (PEFT) has become a popular technique for adapting pre-trained vision foundation models (VFMs) to downstream tasks. Low-ranked adaptation (LoRA) [29] is a widely-used approach for PEFT. Previous works [37, 48] employing LoRA in the VFMs have achieved good empirical generalization results on both deepfake detection and synthetic image detection benchmarks. Compared to full parameter fine-tuning (FFT), LoRA also has the advantage of retaining part of the original well-learned pre-trained semantic knowledge while learning the forgery patterns. Additionally, [38] and [71] introduce MoE-LoRA techniques and dual adapters into the deepfake detection fields. However, our SVD-based approach differs from these adapter-based methods and shows several advantages over them. Specifically, our method explicitly constructs two orthognoal subspaces for semantic and forgery using SVD, ensuring the pre-existing semantic knowledge will not be distorted and well-retained during the process of learning forgery. In contrast, the mentioned adapter-based methods do not explicitly ensure the orthogonality between semantic and forgery, still having the potential to distort the pre-existing semantic knowledge and result in unexpected generalization results.
To show our approach can achieve better generalization results than other adapter-based methods in both deepfake detection and synthetic image detection benchmarks, we provide several empirical results in Tab. 9 and Tab. 10. From these results, we can see that our proposed SVD-based method achieves clearly higher generalization results than adapter-based methods in both deepfake detection and synthetic image detection fields, as our approach explicitly preserves the semantic knowledge while learning the forgery patterns.
Methods FF-c23 FF-c40 Avg.
CNN-Aug [86] 0.8493 0.7846 0.8170
F3Net [64] 0.9635 0.8271 0.8953
SPSL [47] 0.9610 0.8174 0.8892
SRM [53] 0.9576 0.8114 0.8845
RECCE [4] 0.9621 0.8190 0.8906
[…]
Ours 0.9872 0.8429 0.9151
Table 11: Within-domain evaluations using the video-level AUC metric. All detectors are trained on FF-c23 and evaluated on other data. $\dagger$ donates our reproduction using the pre-trained weights, otherwise, we obtain the results from [92].
@@ B.4 Comparison with Existing Adapter-Based Detectors
EfficientNet [80] 82.40 91.11 63.32 80.10
MLGD [41] 84.21 91.82 63.46 77.10
LTW [74] 85.60 92.70 65.60 80.20
DCL [75] 87.70 94.90 68.40 82.93
M2TR [84] 81.07 94.91 55.71 76.99
F3Net [64] 83.57 94.95 61.07 81.20
F2Trans [55] 89.64 97.47 81.43 90.55
CFM [52] 85.00 92.74 76.07 84.55
FA-ViT [51] 92.86 98.10 82.57 91.20
@@
Table 12: Additional results of cross-manipulation evaluation on FF++ (c23). Following [55, 51], we conduct evaluations by training on the other three manipulated methods while testing on the remaining one. Specifically, GID-DF means training on the other three manipulated methods (FF-F2F, FF-FS, FF-NT) while testing on the FF-DF. Results of other methods are cited from [55, 51].
B.5 Robustness Evaluation
To evaluate our model’s robustness to random perturbations, we adopt the methodology used in previous studies [25, 100], which involves examining three distinct types of degradation: Block-wise distortion, Change contrast, and JPEG compression. We apply each of these perturbations at five different levels to assess the model’s robustness under varying degrees of distortion, following [9, 93]. The video-level AUC results for these unseen perturbations, using the model trained on FF++ (c23), are depicted in Fig. 8. Generally, our approach shows higher results than other methods, demonstrating the better robustness of our approach than other models.
Figure 8: Robustness to unseen perturbations. We present video-level AUC for five distinct degradation levels across three types of perturbations in [32].
Appendix C Additional cross-method evaluation
The capability of face forgery detectors to generalize to new manipulation methods is crucial in practical, real-world applications. In our manuscript, we present cross-method evaluations using the DF40 dataset [95]. Specifically, we train the models with four manipulation methods from FF++ (c23) and then test them on the other eight manipulation techniques provided in DF40. Furthermore, we conduct an additional cross-method evaluation following the protocol introduced in [75, 55, 51]. This protocol involves training the model on diverse manipulation types of samples and subsequently testing it on unknown manipulation methods. The results of this evaluation are reported in Tab. 12. It is evident that our proposed method attains remarkable performance in cross-manipulation evaluation. In terms of accuracy (ACC), it outperforms the latest SOTA detector FA-ViT by 2.85% on GID-DF and 3.14% on GID-F2F, respectively.
Figure 9: Logits and confidence distributions of the vanilla CNN (Xception) and ours. We show that (a) the Xception fits the fake patterns in the training set well (the left column), while other information is not adequately captured, leading to poor generalization when detecting unseen deepfake datasets (the right column). In contrast, our approach (b) learns both forgery patterns and semantic knowledge well, achieving more robust generalization performance.
@@ Appendix C Additional cross-method evaluation
C.1 Additional Ablation Studies
Impact of Different Vision Foundation Models
We initialize the ViT backbone with several widely used pre-trained weights from different vision foundation models, including BEIT-v2 [63], CLIP [65], and SigLIP [96]. The results are shown in Tab. 13 and Tab. 14. It is evident that our proposed approach improves the generalization performance of different pre-trained ViTs. On the other hand, we note that different initialization significantly impacts generalization performance, indicating the importance of choosing a suitable pre-trained initialization. Through empirical results, we discover that the ViT pre-trained on CLIP exhibits the highest performance in both deepfake detection and synthetic image detection tasks. Therefore, we choose CLIP as the default setting for our approach.
@@ C.1 Additional Ablation Studies
Table 15: Ablation studies on deepfake image detection (Cross-dataset) regarding different ViT architectures were used. We employ the two architectures implemented in the original paper of CLIP [65] for experiments. All models are trained on FF++ (c23).

### v2

@@ 5 Conclusion
In this paper, we start our research from a new perspective to excavate the failure reason of the generalization in AIGI detection, namely the asymmetry phenomena, where a naively trained detector very quickly shortcuts to the seen fake patterns, collapsing the feature space into a low-ranked structure that limits expressivity and generalization. To address this, we propose integrating higher-ranked pre-trained knowledge from vision foundation models to expand the feature space. Simultaneously, we decompose the feature space into two orthogonal subspaces, for preserving pre-trained knowledge while learning forgery. Beyond LoRA and full-parameters tuning, we explicitly ensure the orthogonality maintaining the higher rank of the whole feature space for better generalization. We conduct extensive experiments with deep analysis on both deepfake and synthetic image detection benchmarks, demonstrating the superior advantages in generalization.
6 Impact Statements
This research advances application-driven machine learning by proposing a novel method for detecting AI-generated images. By effectively identifying deepfakes and curbing malicious uses of generative models, it holds significant potential for positive social impact. However, a possible negative outcome is the misuse of our method to enhance the realism of deepfake generators. To address this concern, we plan to implement measures like access control. Overall, we urge the research community to minimize negative impacts while leveraging the positive contributions of this work.
Appendix A Proof of Theorem-3.2
Proof.
The covariance rank disparity emerges from distinct generation mechanisms. For real data covariance $\Sigma_{\mathrm{real}}$, the factor model assumption implies that $r_{\mathrm{real}}$ independent latent factors with non-degenerate interactions induce $\mathrm{rank}(\Sigma_{\mathrm{real}})\geq r_{\mathrm{real}}$ through singular value decomposition of the factor loading matrix.
For synthetic data, we decompose $\Sigma_{\mathrm{fake}}=\sum_{j=1}^{k}\lambda_{j}\Sigma_{j}+\Sigma_{\epsilon}$ where each component covariance $\Sigma_{j}=\mathbb{E}[G_{j}(z)G_{j}(z)^{\top}]$ satisfies $\mathrm{rank}(\Sigma_{j})\leq r_{\mathrm{fake}}$, and $\Sigma_{\epsilon}$ captures residual covariance. Through rank subadditivity, we obtain $\mathrm{rank}(\Sigma_{\mathrm{fake}})\leq\sum_{j=1}^{k}\mathrm{rank}(\Sigma_{j})+\mathrm{rank}(\Sigma_{\epsilon})\leq kr_{\mathrm{fake}}+d_{\epsilon}$.
Under the Lipschitz continuity condition $\|G_{j}(z)\|_{2}\leq L\|z\|_{2}$ for generators, we apply Gaussian chaos concentration bounds to control the residual term: $\mathbb{P}(\|\Sigma_{\epsilon}\|_{*}>C\sqrt{d\log(2k/\delta)})\leq\delta$. The union bound over $k$ generators consequently establishes the probabilistic rank bound, completing the proof. ∎
Appendix B Algorithm Procedure Illustration
@@
Appendix C Additional Results and Ablations
@@
Table 8: Benchmarking results of cross-method evaluations in terms of Acc performance on the Genimage dataset. We follow (Zhu et al., 2024) and use the SDv1.4 as the training set while others as the testing sets. We directly cite the results of ResNet-50, DeiT-S, Swin-T, CNNSpot, Spec, F3Net, and GramNet from (Zhu et al., 2024). We obtain the results of UnivFD and DRCT from (Chen et al., 2024), and FreqNet, NPR, and FatFormer by using the official checkpoints for reproduction. We report the Accuracy metric for comparison following (Chen et al., 2024).
ResNet-50 (He et al., 2016) CVPR 2016 54.9 99.9 99.7 53.5 61.9 98.2 56.6 52.0 72.1
DeiT-S (Touvron et al., 2021) ICML 2021 55.6 99.9 99.8 49.8 58.1 98.9 56.9 53.5 71.6
Swin-T (Liu et al., 2021b) ICCV 2021 62.1 99.9 99.8 49.8 67.6 99.1 62.3 57.6 74.8
CNNSpot (Wang et al., 2020b) CVPR 2020 52.8 96.3 95.9 50.1 39.8 78.6 53.4 46.8 64.2
Spec (Zhang et al., 2019) WIFS 2019 52.0 99.4 99.2 49.7 49.8 94.8 55.6 49.8 68.8
F3Net (Qian et al., 2020) ECCV 2020 50.1 99.9 99.9 49.9 50.0 99.9 49.9 49.9 68.7
GramNet (Liu et al., 2020) CVPR 2020 54.2 99.2 99.1 50.3 54.6 98.9 50.8 51.7 69.9
UnivFD (Ojha et al., 2023) CVPR 2023 91.5 96.4 96.1 58.1 73.4 94.5 67.8 57.7 79.5
NPR  (Tan et al., 2024c) CVPR 2024 81.0 98.2 97.9 76.9 89.8 96.9 84.1 84.2 88.6
FreqNet  (Tan et al., 2024b) AAAI 2024 89.6 98.8 98.6 66.8 86.5 97.3 75.8 81.4 86.8
FatFormer (Liu et al., 2024) CVPR 2024 92.7 100.0 99.9 75.9 88.0 99.9 98.8 55.8 88.9
DRCT (Chen et al., 2024) ICML 2024 91.5 95.0 94.4 79.4 89.2 94.7 90.0 81.7 89.5
Ours 82.4 99.8 99.8 78.7 93.3 97.4 91.7 77.6 91.1
C.1 Results on GenImage Benchmark
In our manuscript, we present the benchmarking outcomes of the UniversalFakeDetect Dataset. Additionally, we report the results obtained from another widely utilized benchmark known as GenImage (Zhu et al., 2024). This GenImage dataset predominantly utilizes the Diffusion model for image generation, incorporating models such as Midjourney (MidJourney., ), SDv1.4 (Rombach et al., 2022b), SDv1.5 (Rombach et al., 2022b), ADM (Dhariwal et al., 2021), GLIDE (Nichol et al., 2021), Wukong (Wuk, 2022. 5), VQDM (Gu et al., 2022a), and BigGAN (Brock et al., 2018b). Following the settings defined for GenImage, we designate SDv1.4 as the training set and the remaining models as the test set. Given the diverse image sizes within the GenImage dataset, images with a size smaller than 224 pixels are duplicated and subsequently cropped to 224 pixels, following (Tan et al., 2024c). We employ the same setting to re-implement FreqNet, FatFormer, and NPR, and also report the results of UnivFD and DRCT from (Chen et al., 2024).
@@ C.1 Results on GenImage Benchmark
LipForensics (Haliassos et al., 2021) CVPR 2021 0.824 0.735 0.780
FTCN (Zheng et al., 2021) ICCV 2021 0.869 0.740 0.805
HCIL (Gu et al., 2022b) ECCV 2022 0.790 0.692 0.741
RealForensics (Haliassos et al., 2022) CVPR 2022 0.857 0.759 0.808
LTTD (Guan et al., 2022) NeurIPS 2022 0.893 0.804 0.849
AltFreezing (Wang et al., 2023b) CVPR 2023 0.895
TALL-Swin (Xu et al., 2023) ICCV 2023 0.908 0.768 0.838
StyleDFD (Choi et al., 2024) CVPR 2024 0.890
NACO (Zhang et al., 2024) ECCV 2024 0.895 0.767 0.831
@@
C.2 Comparison with Existing Video Detectors
@@
In the manuscript, we mainly compare our method with image detectors. Here, we provide an individual result to compare our approach with existing SOTA video detectors. Following (Zhang et al., 2024; Xu et al., 2023), we conduct evaluations on the widely-used CDF-v2 (Li et al., 2020b) and DFDC (detection challenge., 2020) using the video-level AUC metric. We have considered both the classical detectors such as LipForensics and the latest SOTA detectors such as NACO (ECCV’24) for a comprehensive comparison. Results in Tab. 9 demonstrate that our image-based approach achieves higher generalization performance in both CDF-v2 and DFDC, improving 4.5% and 3.9% points than the second-best video-based models. This further validates the superior generalization performance of our approach.
Table 10: Additional results of cross-manipulation evaluation on FF++ (c23). Following (Miao et al., 2023; Luo et al., 2023a), we conduct evaluations by training on the other three manipulated methods while testing on the remaining one. Specifically, GID-DF means training on the other three manipulated methods (FF-F2F, FF-FS, FF-NT) while testing on the FF-DF. Results of other methods are cited from (Miao et al., 2023; Luo et al., 2023a).
@@
EfficientNet (Tan & Le, 2019) 82.40 91.11 63.32 80.10
MLGD (Li et al., 2018) 84.21 91.82 63.46 77.10
LTW (Sun et al., 2021) 85.60 92.70 65.60 80.20
DCL (Sun et al., 2022) 87.70 94.90 68.40 82.93
M2TR (Wang et al., 2022) 81.07 94.91 55.71 76.99
F3Net (Qian et al., 2020) 83.57 94.95 61.07 81.20
F2Trans (Miao et al., 2023) 89.64 97.47 81.43 90.55
CFM (Luo et al., 2023b) 85.00 92.74 76.07 84.55
FA-ViT (Luo et al., 2023a) 92.86 98.10 82.57 91.20
@@
C.3 Robustness Evaluation
To evaluate our model’s robustness to random perturbations, we adopt the methodology used in previous studies (Haliassos et al., 2021; Zheng et al., 2021), which involves examining three distinct types of degradation: Block-wise distortion, Change contrast, and JPEG compression. We apply each of these perturbations at five different levels to assess the model’s robustness under varying degrees of distortion, following (Chen et al., 2022b; Yan et al., 2024a). The video-level AUC results for these unseen perturbations, using the model trained on FF++ (c23), are depicted in Fig. 7. Generally, our approach shows higher results than other methods, demonstrating the better robustness of our approach than other models.
Figure 7: Robustness to unseen perturbations. We present video-level AUC for five distinct degradation levels across three types of perturbations in (Jiang et al., 2020).
@@ C.3 Robustness Evaluation
Appendix D Additional cross-method evaluation
The capability of face forgery detectors to generalize to new manipulation methods is crucial in practical, real-world applications. In our manuscript, we present cross-method evaluations using the DF40 dataset (Yan et al., 2024b). Specifically, we train the models with four manipulation methods from FF++ (c23) and then test them on the other eight manipulation techniques provided in DF40. Furthermore, we conduct an additional cross-method evaluation following the protocol introduced in (Sun et al., 2022; Miao et al., 2023; Luo et al., 2023a). This protocol involves training the model on diverse manipulation types of samples and subsequently testing it on unknown manipulation methods. The results of this evaluation are reported in Tab. 10. It is evident that our proposed method attains remarkable performance in cross-manipulation evaluation. In terms of accuracy (ACC), it outperforms the latest SOTA detector FA-ViT by 2.85% on GID-DF and 3.14% on GID-F2F, respectively.
Table 13: Ablation studies on deepfake image detection (Cross-dataset) regarding different ViT architectures were used. We employ the two architectures implemented in the original paper of CLIP (Radford et al., 2021) for experiments. All models are trained on FF++ (c23).
@@ Appendix D Additional cross-method evaluation
Figure 8: t-SNE visualizations of the initial latent feature spaces between vanilla CNN (Wang et al., 2020b) and CLIP (Radford et al., 2021). We show that both the pre-trained CNN and CLIP can identify different semantic objects.
Figure 9: t-SNE visualizations of the latent feature spaces between vanilla CNN (Wang et al., 2020b) and ours. We use the testing set of ProGAN and StyleGAN within UniversalFakeDetect Dataset (Wang et al., 2020b) for visualization. We see that after fine-tuning on AIGI data, the baseline quickly shortcuts to the fake and “forgets” the pre-existing semantic knowledge, thereby resulting in a highly constrained feature space.
Figure 10: t-SNE visualizations of the latent feature spaces between vanilla CNN (Wang et al., 2020b) and ours. We use the testing set of CycleGAN within UniversalFakeDetect Dataset (Wang et al., 2020b) for visualization.
D.1 Additional Ablation Studies
Impact of Different Vision Foundation Models
We initialize the ViT backbone with several widely used pre-trained weights from different vision foundation models, including BEIT-v2 (Peng et al., 2022), CLIP (Radford et al., 2021), and SigLIP (Zhai et al., 2023). The results are shown in Tab. 11 and Tab. 12. It is evident that our proposed approach improves the generalization performance of different pre-trained ViTs. On the other hand, we note that different initialization significantly impacts generalization performance, indicating the importance of choosing a suitable pre-trained initialization. Through empirical results, we discover that the ViT pre-trained on CLIP exhibits the highest performance in both deepfake detection and synthetic image detection tasks. Therefore, we choose CLIP as the default setting for our approach.